OpenStudy (anonymous):

write the polynomial equation whose roots (zeros) are: 2, -1, 5

OpenStudy (anonymous):

what is the comma doing there?

OpenStudy (anonymous):

the first statment is the question. the statement after that i was wondering if that was the answer to the question.. if that makes any sense

OpenStudy (anonymous):

oh i see your answer should be a number \[3x^2=12x-15\] \[3x^2-12x+15=0\]divide by 3 get \[x^2-4x+5=0\] then use the quadratic formula since this one will not factor using integers

OpenStudy (anonymous):

in fact there is no real solution

OpenStudy (anonymous):

hold the phone

OpenStudy (anonymous):

this line 3x(x-5)-3(x-5)=0 is incorrect

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

check it is not what you started with

OpenStudy (anonymous):

i have to solve by compeleting the square.. prob should have told yall that earlier

OpenStudy (anonymous):

\[3x^2 -15x + 3x +15\neq 3x(x-5)-3(x-5)=0\]

OpenStudy (anonymous):

it has no real solution, but we can find the complex ones and we can complete the square easily

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

and therefore \[x=2\pm\sqrt{11}i\] but i am going to guess that they is a typo in the problem look carefully and make sure that the problem you posted is the one you were given

OpenStudy (radar):

Did you divide the -15 by 3?

OpenStudy (anonymous):

oh lord a sophomore mistake

OpenStudy (anonymous):

yes! the problem is 3x^2=12x-15

OpenStudy (radar):

Jus t watchin.

OpenStudy (anonymous):

\[x^2-4x=-5\] that is better then \[(x-2)^2=-5+4=-1\] \[x-2=\pm\sqrt{-1}=\pm i\] \[x=2\pm i\]

OpenStudy (anonymous):

i was so pleased with myself for dividing by 3 the first time i forgot the second time. sorry

OpenStudy (radar):

First time was better.

OpenStudy (anonymous):

dont you have to move the -5 over to the otherside of the equation and make it a positive 5

OpenStudy (anonymous):

not if you want to complete the square, no idea is to make it look like \[(x+b)^2=a\] and then take the square root so best to leave the constant on the other side

OpenStudy (anonymous):

thank you!!!!!!!! more problems though :) how do i graph the following f(x)=lx+3l

OpenStudy (anonymous):

now teach me dont give me the answers if you can. i need to learn if ya have time

OpenStudy (radar):

Do you know what "absolute value" means?

OpenStudy (anonymous):

yes, with a calculator..

OpenStudy (radar):

What you have is a absolute function f(x)=|x+3| the vertical bars on each side of the x=3 indicate the absolute value. So you could look at that like this: y= +(x+3) and y= -(x+3)

OpenStudy (radar):

Now simply graph the following: y=x+3 and get one line and on the same graph, graph the other value y=-x-3. this should be the same as x+3=y and x+3=-y.

OpenStudy (radar):

There is a typo in the second post above *x=3, x+3

OpenStudy (anonymous):

so i just draw a graph??

OpenStudy (radar):

 something like that